# Non-Contact Health Monitoring During Daily Personal Care Routines

### v1

@@ Abstract
Remote photoplethysmography (rPPG) enables non-contact, continuous monitoring of physiological signals and offers a practical alternative to traditional health sensing methods. Although rPPG is promising for daily health monitoring, its application in long-term personal care scenarios—such as mirror-facing routines in high-altitude environments—remains challenging due to ambient lighting variations, frequent occlusions from hand movements, and dynamic facial postures. To address these challenges, we present LADH (Long-term Altitude Daily Health), the first long-term rPPG dataset containing 240 synchronized RGB and infrared (IR) facial videos from 21 participants across five common personal care scenarios, along with ground-truth PPG, respiration, and blood oxygen signals. Our experiments demonstrate that combining RGB and IR video inputs improves the accuracy and robustness of non-contact physiological monitoring, achieving a mean absolute error (MAE) of 4.99 BPM in heart rate estimation. Furthermore, we find that multi-task learning enhances performance across multiple physiological indicators simultaneously. Dataset and code are open at https://github.com/McJackTang/FusionVitals.
Index Terms:
non-contact, rPPG, health-monitoring
@@ I Introduction
To address these challenges, we present LADH, a multi-modal biosensing dataset for daily personal care activities. LADH contains 240 synchronized RGB-IR facial videos collected across five scenarios: sitting at rest, sitting during personal care (toothbrushing/hair-combing with facial occlusions), standing at rest, standing during personal care, and post-exercise. Eleven participants contributed data over a continuous 10-day period, with all recordings synchronized at millisecond precision with ground-truth physiological signals from pulse oximeters and respiratory bands [7].
Our evaluation using both subject-independent and day-wise partitioning reveals that day-wise partitioning enhances measurement accuracy, while RGB-IR video fusion with multi-task learning significantly improves physiological signal prediction despite frequent facial occlusions.
@@
  • The first long-term rPPG dataset featuring synchronized RGB-IR videos with ground-truth physiological measurements across daily personal care scenarios with realistic facial occlusions.
@@
SUMS [7] 80 Face+Finger PPG/SpO2/RR
LADH 240 Face(RGB+IR) PPG/SpO2/RR
@@ II-A Physiological Sensing in Daily Personal Care Scenarios
Previous studies have demonstrated the feasibility of smart mirror systems for non-contact, real-time heart rate measurement without external sensors [11, 12]. These results indicate that non-contact health monitoring systems can offer a passive, user-friendly, and privacy-preserving approach in daily personal care scenarios. However, there is limited research on the robustness of such systems in long-term, comprehensive user studies. The LADH dataset aims to address this gap by focusing on daily personal care routines in a real-world context.
@@ III Dataset
Approved by the Institutional Review Board, we recruited 21 graduate students (aged 23–28) as participants and conducted data collection in a simulated daily personal care scenario. The WN-12207K3321SM290 camera module was used to capture participants’ facial videos in both RGB and IR modalities. Concurrently, physiological ground-truth signals were recorded using a CMS50E pulse oximeter for photoplethysmography (PPG) and SpO2, and an HKH-11C respiratory sensor to monitor breathing patterns. Video recordings were acquired at a resolution of 640×480 pixels and a frame rate of 30 frames per second (FPS). PPG signals were sampled at 20 Hz, while RR were recorded at 50 Hz. All video frames and physiological signals were synchronously collected at millisecond-level precision using a data acquisition platform based on the PhysRecorder framework [13].
@@ III-A Data Collection
The data collection protocol is illustrated in Figure 1. We organized participants into two groups: 10 subjects who completed five scenarios in a single day, and 11 subjects who performed the same scenarios daily for 10 consecutive days. For the seated resting condition (state 1), participants wore an HKH-11C respiratory sensor on their abdomen and a CMS50E pulse oximeter on their left index finger while sitting upright facing the camera. Participants maintained minimal movement with a fixed gaze at the camera during the two-minute recording period.
In state 2, participants remained seated with the same equipment while performing toothbrushing and hair-combing activities for two minutes. State 3 involved a standing resting position with the same sensor configuration, also recorded for two minutes. In state 4, participants repeated the toothbrushing and hair-combing actions while standing for two minutes. After completing these four conditions, participants engaged in moderate physical exercise (squats, high knees, or breath-holding) to induce physiological changes. Immediately following exercise, participants were recorded for an additional two minutes in a seated position (state 5). The exercise component was included to produce variations in SpO2 levels. The experimental setup during data collection is shown in Figure 2.
@@ IV-A Input embedding
Our FusionPhysNet introduces an innovative input embedding strategy, demonstrating the model’s ability to process video signals from two distinct modalities: RGB video facial and IR video facial. As illustrated in Figure 3, the model first performs pre-processing on both types of video data to convert them into frame sequences. These sequences are then passed through a shared 3D convolutional encoder for spatiotemporal feature extraction, followed by a dedicated fusion module (FusionNet) that integrates the extracted features.
This architecture enables the model to jointly exploit complementary information from both RGB and IR facial videos. By integrating features from both modalities, the model achieves a more comprehensive understanding of an individual’s physiological state, effectively combining superficial and deeper physiological signals. This multimodal fusion framework significantly enhances the model’s performance in complex real-world scenarios and establishes a more accurate and reliable foundation for vital sign estimation, including heart rate and blood oxygen saturation.
@@ IV-B Neural Network Model
The overall architecture of the FusionPhysNet is illustrated in Figure 3. We introduce a novel design in the FusionNet module by incorporating a modality-aware fusion mechanism. Specifically, a gated feature selection strategy is employed to adaptively modulate the contribution of each modality based on its global contextual representation, thereby effectively integrating information from both facial RGB and facial IR video streams. This design enables the model to dynamically emphasize the more informative modality under varying environmental conditions (e.g., changes in illumination), significantly enhancing the robustness and generalizability of the physiological signal estimation framework.
TABLE II: Results of HR-SpO2-RR Multi-Task Training by Subject
HR Task SpO2 Task RR Task
@@
IV-C Joint-Training
Loss $=\mathrm{MSE}_{\mathrm{BVP}}+\mathrm{MSE}_{\mathrm{RR}}$
[…]
V Results and Findings
We conducted two sets of experiments based on different dataset partitioning strategies: subject-wise partitioning and day-wise partitioning. We established different training tasks, including individual training for HR, SpO2 and RR under multimodal input, joint training for HR-SpO2-RR under multimodal input, joint training for HR-SpO2-RR under RGB-only video input and IR-only video input.
@@ V Results and Findings
TABLE III: Results of HR-SpO2-RR Multi-Task Training by Day
HR Task SpO2 Task RR Task
@@ VI Limitations
While this study provides valuable insights into the feasibility of non-contact health monitoring in daily personal care scenarios, it has certain limitations. The relatively small sample size may not fully capture the diversity of physiological characteristics and behavioral patterns across the general population. In addition, due to the limitations of the dataset, we were unable to conduct certain comparative experiments to further validate the capabilities of our model. Future research should aim to include a larger and more diverse participant pool and incorporate a broader range of real-world personal care behaviors to enhance the generalizability and robustness of the findings.
VII Conclusions
This study proposes a non-contact physiological monitoring framework using facial RGB and IR videos in daily personal care scenarios. The LADH dataset demonstrates that modality fusion with multi-task learning significantly improves performance. In cross-subject settings, HR MAE dropped from 9.02 to 7.12 and RR from 2.25 to 1.43; in cross-day settings, HR decreased from 5.23 to 4.99. These gains reveal the complementary value of RGB and IR signals and the effectiveness of joint HR–SpO2–RR prediction. Multi-task learning captures shared physiological patterns, especially for periodic signals, enhancing generalization. This work lays a foundation for accurate, roabut snd non-invasive health monitoring in everyday scenarios.

### v2

@@ Abstract
Remote photoplethysmography (rPPG) enables non-contact, continuous monitoring of physiological signals and offers a practical alternative to traditional health sensing methods. Although rPPG is promising for daily health monitoring, its application in long-term personal care scenarios—such as mirror-facing routines in high-altitude environments—remains challenging due to ambient lighting variations, frequent occlusions from hand movements, and dynamic facial postures. To address these challenges, we present the Long-term Altitude Daily Health (LADH) dataset, the first long-term rPPG dataset containing 240 synchronized RGB and infrared (IR) facial videos from 21 participants across five common personal care scenarios, along with ground-truth PPG, respiration, and blood oxygen signals. Our experiments demonstrate that combining RGB and IR video inputs improves the accuracy and robustness of non-contact physiological monitoring, achieving a mean absolute error (MAE) of 4.99 BPM in heart rate estimation. Furthermore, we find that multi-task learning enhances performance across multiple physiological indicators simultaneously. Dataset and code are open at https://github.com/McJackTang/FusionVitals.
@@ I Introduction
To address these challenges, we present Long-term Altitude Daily Health (LADH) dataset, a multi-modal biosensing dataset for daily personal care activities. LADH contains 240 synchronized RGB and IR facial videos collected across five scenarios: sitting at rest, sitting during personal care (toothbrushing/hair-combing with facial occlusions), standing at rest, standing during personal care, and post-exercise. Eleven participants contributed data over a continuous 10-day period, with all recordings synchronized at millisecond precision with ground-truth physiological signals from pulse oximeters and respiratory bands [7].
Our evaluation using both subject-independent and day-wise partitioning reveals that day-wise partitioning enhances measurement accuracy, while RGB and IR video fusion with multi-task learning significantly improves physiological signal prediction despite frequent facial occlusions.
@@
  • The first long-term rPPG dataset featuring synchronized RGB and IR videos with ground-truth physiological measurements across daily personal care scenarios with realistic facial occlusions.
@@
SUMS [7] 80 Face+Finger PPG/SpO2/$R$
LADH 240 Face(RGB+IR) PPG/SpO2/$R$
@@ II-A Physiological Sensing in Daily Personal Care Scenarios
Previous studies have demonstrated the feasibility of smart mirror systems for non-contact, real-time heart rate measurement without external sensors [11, 12]. These results indicate that non-contact health monitoring systems can offer a passive, user-friendly, and privacy-preserving approach in daily personal care scenarios. However, there is limited research on the robustness of such systems in long-term, comprehensive user studies. The LADH dataset aims to address this gap by focusing on daily personal care routines in the real world.
@@ III Dataset
Approved by the Institutional Review Board, we recruited 21 graduate students (aged 23–28) as participants and conducted data collection in a simulated daily personal care scenario. The WN-12207K3321SM290 camera module was used to capture participants’ facial videos in both RGB and IR modalities. As shown in Figure 1, physiological ground-truth signals were recorded using a CMS50E pulse oximeter for photoplethysmography (PPG) and SpO2, and an HKH-11C respiratory sensor to monitor breathing patterns. Video recordings were acquired at a resolution of 640×480 pixels and a frame rate of 30 frames per second (FPS). PPG signals were sampled at 20 Hz, while RR signals were recorded at 50 Hz. All video frames and physiological signals were synchronously collected at millisecond-level precision using a data acquisition platform based on the PhysRecorder [13].
@@ III-A Data Collection
The data collection protocol is illustrated in Figure 2. Data were collected from 21 participants in five scenarios, where 11 subjects participated in a 10-day longitudinal study, while the remaining 10 contributed single-day recordings. For the seated resting condition (state 1), participants wore an HKH-11C respiratory sensor on their abdomen and a CMS50E pulse oximeter on their left index finger while sitting upright facing the camera. Participants maintained minimal movement with a fixed gaze at the camera during the two-minute recording period.
In state 2, participants remained seated with the same equipment while performing toothbrushing and hair-combing activities for two minutes. State 3 involved a standing resting position with the same sensor configuration, also recorded for two minutes. In state 4, participants repeated the toothbrushing and hair-combing actions while standing for two minutes. After completing these four conditions, participants engaged in moderate physical exercise (squats, high knees, or breath-holding) to induce physiological changes. Immediately following exercise, participants were recorded for an additional two minutes in a seated position (state 5). The exercise component was included to produce variations in SpO2 levels.
@@ IV-A Input embedding
Building upon PhysNet [14, 7], our FusionPhysNet introduces an input embedding strategy that processes facial video from two modalities: RGB and IR. As shown in Figure 3, both video streams are pre-processed into frame sequences and passed through a shared 3D convolutional encoder for spatiotemporal feature extraction. By integrating features from both modalities, the model achieves a more comprehensive understanding of an individual’s physiological state, effectively combining superficial and deeper physiological signals. This multimodal fusion framework significantly enhances the model’s performance in complex real-world scenarios and establishes a more accurate and reliable foundation for vital sign estimation, including heart rate and blood oxygen saturation.
@@ IV-B Neural Network Model
The overall architecture of the FusionPhysNet is illustrated in Figure 3. We extend the PhysNet [14] backbone with a modality-aware fusion mechanism in FusionNet. Specifically, a gated feature selection strategy adaptively modulates the contribution of each modality based on global contextual representations. This design enables the model to dynamically emphasize the more informative modality under varying environmental conditions (e.g., changes in illumination), enhancing the robustness and generalizability of the physiological signal estimation framework.
TABLE II: Intra-Dataset and Inter-Dataset Experiment
Training Set LADH SUMS PURE
[…]
LADH 8.15 9.19 16.93 18.20 17.00 18.78
SUMS 11.23 15.45 3.36 3.84 14.95 17.11
PURE 8.10 8.83 7.97 8.87 0.59 0.77
The table shows the cross-dataset experimental results of the LADH, SUMS [7], and PURE [5] datasets on the PhysNet [14] model.
IV-C Joint-Training
Loss $=\mathrm{MSE}_{\mathrm{BVP}}+\mathrm{MSE}_{\mathrm{RR}}$
where $\text{MSE}_{\text{BVP}}$ represents the mean squared error of the blood volume pulse (BVP) signal, $\text{MSE}_{\text{RR}}$ represents the mean squared error of the RR signal, and $\text{MSE}_{\text{SpO${}_{2}$}}$ represents the mean squared error of the SpO2 signal. This formulation aims to balance the precision of the BVP, RR and SpO2 measurements, leading to a more robust model for multimodal physiological signal measurement.
V Results and Findings
We validated the effectiveness of LADH through both intra- and inter-dataset experiments using PhysNet [14]. As shown in Table II, LADH achieves reasonable intra-dataset accuracy (MAE=8.15 BPM) and maintains competitive cross-dataset performance when tested on SUMS and PURE, outperforming SUMS in cross-transfer to PURE (MAE 8.10 vs. 14.95). These results highlight LADH’s dual role: it provides a strong baseline within its own domain while also exposing the inherent challenges of generalization across different datasets, underscoring its value as a benchmark for robust rPPG evaluation.
We then conducted two sets of experiments based on different dataset partitioning strategies within LADH: subject-wise partitioning and day-wise partitioning. We established different training tasks, including individual training for HR, SpO2 and RR under multimodal input, joint training for HR-SpO2-RR under multimodal input, joint training for HR-SpO2-RR under RGB-only video input and IR-only video input.
TABLE III: Results of HR-SpO2-$R$ Multi-Task Training by Subject
Modality HR Task SpO2 Task RR Task
@@ V Results and Findings
TABLE IV: Results of HR-SpO2-$R$ Multi-Task Training by Day
Modality HR Task SpO2 Task RR Task
@@ VI Limitations
While this study provides valuable insights into the feasibility of non-contact health monitoring in daily personal care scenarios, it has certain limitations. The relatively small sample size and age range may not fully capture the diversity of physiological characteristics and behavioral patterns across the general population. Future research should aim to include a larger and more diverse participant pool and incorporate a broader range of real-world personal care behaviors to enhance the generalizability and robustness of the findings.
VII Conclusion
This study introduces the LADH dataset, the first long-term rPPG dataset with synchronized RGB, IR, and ground-truth physiological signals across realistic daily care scenarios. By releasing 240 multimodal videos with PPG, respiration, and SpO2, LADH provides a valuable benchmark for robust non-contact health monitoring. We further propose FusionPhys, a modality-fusion and multi-task learning framework that improves accuracy and generalization, reducing HR and RR errors across both cross-subject and cross-day settings. FusionPhys reveals the complementary value of RGB and IR signals and the effectiveness of joint HR–SpO2–RR prediction. This work lays a foundation for accurate, robust, and non-invasive health monitoring in everyday scenarios.
Acknowledgment
This work is supported by the foundation of National Key Laboratory of Human Factors Engineering No. HFNKL2024W06, the National Natural Science Foundation of China No. 62366043 & 62472244, Beijing Natural Science Foundation No.QY24248, the National Key R&D Program of China No. 2024YFB4505500, No. 2024YFB4505503, the Tsinghua University Initiative Scientific Research Program No. 20257020004, and Qinghai University Research Ability Enhancement Project No. 2025KTSA05.